OpenStudy (anonymous):

Can anyone help me with my kkt problem? MAX E = -x1logx1 -x2logx2 - x3logx3 st x1 + x2 + x3 = 1 0x1 + .5x2 + x3 = * x1 >= 0, x2 >= 0, x3 >=0 and * has 9 values between .1 and .9

OpenStudy (anonymous):

This is what I have so far: -logx1 -1 - lamda1= 0 -logx2 - 1 - lamda1 -.5lamda2 = 0 -logx3 - 1 - lamda1 - lamda2 = 0 and matrix [1 1 1 0 0] [0 .5 1 0 0] [? 0 0 -1 -1] [0 ? 0 -1 -.5] [0 0 ? -1 -1]